### v3

# GQSA: Group Quantization and Sparsity for Accelerating Large Language Model Inference

Abstract
Model compression has emerged as a mainstream solution to reduce memory usage and computational overhead. This paper presents Group Quantization and Sparse Acceleration (GQSA), a novel compression technique tailored for LLMs. Traditional methods typically focus exclusively on either quantization or sparsification, but relying on a single strategy often results in significant performance loss at high compression rates. In contrast, GQSA integrates quantization and sparsification in a tightly coupled manner, leveraging GPU-friendly structured group sparsity and quantization for efficient acceleration. Building upon system-algorithm co-design principles, we propose a two-stage sparse optimization strategy that ensures the performance superiority of the compressed model. On the engine side, we introduce a "task-centric" parallel strategy, which, to the best of our knowledge, is the first application in the domain of sparse computing. Compared to the traditional 2:4 sparse method, the GQSA offers a more flexible and adjustable sparsity rate, as well as a higher weight compression rate, and is efficiently compatible with weight-only quantization methods. Experimental results demonstrate that, under the GQSA W4S50% compression setting, the model’s accuracy surpasses that of both 2:4 pruning and W2 quantization. Furthermore, at the inference level, GQSA outperforms W2 by 1.26$\times$ and 2:4 pruning by 2.35$\times$ in terms of speed.
GQSA: Group Quantization and Sparsity for Accelerating Large Language Model Inference
1 Introduction
Sparsity, combined with quantizationLin et al. (2024); Shao et al. (2023), is a powerful approach to enhance model inference performance, reduce the size of LLMs, and enable their deployment on edge devices such as PCsGu et al. (2024); Liu et al. (2022). However, current sparsification strategies exhibit limited acceleration benefits due to the unstructured sparsity patterns typically generated by existing unstructured pruning methodsHan et al. (2015); Sun et al. (2023), which are poorly suited for hardware acceleration. Strategies such as SparseGPTFrantar and Alistarh (2023) and WandaSun et al. (2023) address this issue by adopting a 2:4 sparsity pattern, leveraging NVIDIA GPUs’ Sparse Tensor Core units for acceleration. Nevertheless, these approaches are constrained by hardware requirements such as a minimum operation shape of [m, n, k] = [16, 8, 16], which restrict their applicability to compute-intensive tasks like GEMM operations. These limitations pose significant challenges in accelerating the decoding process, the primary performance bottleneck in LLMsZeng et al. (2024). Unlike GEMM, decoding involves GEMV operations, where the Tensor Core’s compute resources are underutilized, with approximately 87.5% of resources being wastedMishra et al. (2021). Consequently, SparseGPT and Wanda achieve up to 50% sparsity but remain inefficient in practical scenarios. Furthermore, these methods are incompatible with weight-only quantization, as Sparse Tensor Cores require both weights and activations to are floating-point or integer formats. Combining sparsification with weight-activation quantization leads to excessive compression of activation value representation, resulting in severe performance degradation. This limitation significantly diminishes the practical utility of existing sparsification strategies.
To address these challenges, we propose a novel model compression method called GQSA, designed specifically for the decoding process and efficiently compatible with weight-only per-group quantization. GQSA explores a group sparsity pattern beyond the conventional 2:4 sparsity, achieving a better trade-off between accuracy and speed through a combination of algorithm-level optimizations and a customized software engine. Specifically, we reinterpret weight pruning as a particular form of quantization and introduce a group pruning based on group quantization. Our method incorporates the Block Sparse Row (BSR) format and designs a compact, low-precision weight storage structure to maximize the compression benefits of pruning and quantization. The GQSA method consists of two main stages. The first stage, Block Quantization-Pruning Optimization (BQPO), calibrates model parameters at the block level by optimizing weight distributions within block to minimize performance loss caused by group pruning and quantization. In the second stage, End-to-End Optimized Quantization-Pruning (E2E-OQP), the backbone network’s weights are frozen, and only the quantization parameters are fine-tuned to optimize the global network performance. Unlike BQPO, E2E-OQP considers the global error distribution across blocks. Freezing the backbone network can not only reduce memory usage but also improve training efficiency. Extensive experiments demonstrate that GQSA achieves significant advantages in both model accuracy and inference speed, especially when applied to newly released advanced models such as LLaMA-3 and LLaMA-3.1 model family.
In summary, our contributions are as follows.
  • We propose a sparse scheme seamlessly compatible with widely used weight-only per-group quantization, effectively accelerating GEMV operations and reducing memory usage.
  • We introduce a task-centric parallel implementation, addressing the workload balancing issue in sparse acceleration.
  • We integrate group pruning with low-bit quantization techniques and achieves outstanding model performance through the two-stage optimization process of BQPO and E2E-OQP.
2 Related work
Compressing Large Language Models. Pruning and quantization are the two primary techniques for compressing LLMs. Pruning methods can be classified into structuredChen et al. (2024); Ma et al. (2023); Ashkboos et al. (2024), semi-structuredciteFrantar and Alistarh (2023); Sun et al. (2023); Fang et al. (2024), and unstructuredHan et al. (2016, 2015); Sun et al. (2023) pruning, depending on the granularity of pruning. Structured pruning operates at a coarser granularity and offers significant acceleration, but it often results in a substantial loss of accuracyWang et al. (2024), limiting its application in LLMs. Unstructured pruning better preserves accuracy but provides limited improvements in inference speedFrantar and Alistarh (2023). Semi-structured pruning strikes a balance between accuracy retention and acceleration, though it is constrained by a sparsity of 50%, reducing its flexibility. Quantization reduces model size by replacing floating-point numbers with low-precision integers, which accelerates memory access during inference. Currently, high-bit quantization techniques such as AWQLin et al. (2024), GPTQFrantar et al. (2022), QuIPChee et al. (2024), OmniQuantShao et al. (2023), and OWQLee et al. (2024) are widely adopted. However, extremely low-bit quantization poses significant challenges, with mainstream methods struggling to maintain performance at low-bit levels. While techniques like AQLMEgiazarian et al. (2024) and QuIP#Tseng et al. (2024) aim to enhance low-bit quantization, they rely on vector quantization and complex codebooks, which hinder inference acceleration. Overall, existing model compression techniques continue to face substantial challenges in achieving an optimal balance between flexibility and compression rate.
Figure 1: The distribution of the top 1% significant weights in the Hessian matrix, derived from the $k_{proj}$ and $q_{proj}$ distributions in the LLaMA-7B model.
Advantages of GQSA. Quantization and sparsity address model redundancy in different ways. Quantization reduces the precision of numerical representations, while sparsity compresses the model by eliminating certain neurons. These two techniques are largely orthogonal, and GQSA leverages both dimensions to achieve flexible and high compression rates. Although both GQSA and 2:4 pruning are semi-structured pruning methods, GQSA offers several advantages over 2:4 pruning. First, GQSA supports an adjustable sparsity rate, whereas 2:4 pruning, designed for NVIDIA’s 2:4 TensorCore, mandates a 50% sparsity rate by forcing two out of every four weights to be zero. Our group sparsity model, in combination with co-designed operators, enables efficient implementation at various sparsity levels. Second, GQSA achieves a higher weight compression rate. For instance, with a 50% sparsity rate, 2:4 pruning requires additional metadata to identify the positions of retained neurons, which are chosen randomly. In contrast, GQSA stores location information at the group/block level, significantly enhancing compression efficiency. Finally, 2:4 pruning is restricted to NVIDIA’s 2:4 TensorCore and is incompatible with mainstream weight-only quantization methods. In contrast, GQSA is highly compatible with weight-only quantization, thanks to its customized two-stage optimization, leading to a substantial increase in overall compression rate.
3 GQSA
3.1 Preliminary
Weight Quantization. LLM quantization maps floating-point values to a lower-bit discrete value space, significantly reducing model size, enhancing computational efficiency, and accelerating inference. The process typically involves two steps: determining the quantization parameters (scale and zero-point) and computing the corresponding quantized tensor. For uniform asymmetric quantization, which is used in this paper, the scale $s$ and zero-point $z$ are determined by:
[…]
where $W$ represents the model weights and $n$ denotes the quantization bit-width. The elements of the quantized tensor can be computed as follows:
$\tilde{W}=\mathrm{clamp}(\left\lfloor\frac{W}{s}\right\rceil+z,0,2^{n}-1),$ (2)
[…]
$\hat{W}=(\tilde{W}-z)\cdot s,$ (3)
where $\hat{W}$ denotes the dequantized weights utilized in the forward computation. The processes of quantization (as shown in Equation (2) and dequantization (as shown in Equation (3) are integrated into the computational graph, enabling quantization-aware optimization through gradient descent.
Salient Weight. In LLMs, different weights exhibit different importance. By pruning unimportant weights, memory usage can be greatly reduced while maintaining nearly unchanged performance. Early studies used the absolute value of weights to evaluate weight importance, but ignored the role of activation. The Hessian metric combines weights and activations and is a more effective metric that has been verified by multiple methodsShang et al. (2023); Frantar and Alistarh (2023). Therefore, this paper uses the Hessian matrix to evaluate weight importance.
$s_{i}=\frac{w_{i}^{2}}{\left[H^{-1}\right]_{ii}^{2}},$ (4)
where $H$ represents the Hessian matrix of each layer, and $w_{i}$ denotes the weight values. In the subsequent sections, $s_{i}$ refers to the criteria for identifying salient weights. AWQLin et al. (2024) demonstrates that the top 1% of salient weights in the model are crucial to performance, so accurately retaining these weights is key to performance. Figure 1 visualizes the distribution of salient weights in the OPT model, revealing a segmented pattern along the rows. Consequently, group by rows and selecting salient weight group emerges as a natural optimization strategy.
Figure 2: GQSA computes saliency metrics based on weights and activations, grouping the weights along the row dimension (illustrated with groups of four elements). Group pruning is then applied based on the average saliency metrics within each group, resulting in the formation of the GQS layer.
Figure 3: Overview of GQSA. (a) We propose a group quantization and sparse LLMs, where linear layers are replaced by GQS layers. (b) We use the two-stage optimization method BQPO and E2E-OQP to recover the performance of the extremely compressed model.
3.2 GQS Layer
Weight-only per-group quantization has gained significant recognition in both academiaLin et al. (2024); Shao et al. (2023); Frantar et al. (2022) and industryGerganov (2024). To enable efficient sparse acceleration compatible with weight-only per-group quantization approach, we conduct a comprehensive analysis of the distribution of salient weights within the model. As depicted in Figure 1, we observe that salient weights exhibit distinct segmented distribution patterns. Based on this observation, we introduce a novel structured group pruning method that goes beyond the conventional 2:4 sparsity pattern, leveraging the segmented distribution characteristics of salient weights. As illustrated in Figure 2, we begin by grouping weights along the row dimension, assuming a group size of 4 for simplicity. For each group, we compute a salient metric using the Hessian matrix. Based on this metric, we prune non-salient weight groups and quantize the remaining salient groups to 4 bits, thereby further compressing the model size. Additionally, by adopting the BSR sparse format, we convert the compression gains from pruning into actual storage savings. The specific storage structure in Figure 2 is shown below:
rowIndex = {0, 1, 3, 3, 4}
groups = {1, 0, 1, 1}
values = {5, 1, 15, 1, 15, 13, 2, 1,
-1, 7, 14, 0, 0, 3, 6, 15}
where rowIndex[i] represent the offset of each row i, where i belongs to the range $[0,\text{rows}]$. The difference rowIndex[r+1] - rowIndex[i] indicates the number of non-zero groups in the i-th row. Additionally, rowIndex[rows] represents the total number of non-zero groups. The array groups[i] stores the indices of the non-zero groups; for instance, if groups[1] = 0, it means that the second group is located in the 0th column (in terms of group units). Finally, values stores the values of the non-zero groups for each row.
3.3 BQPO
In the first stage (shown in Figure 3(b)), we apply the BQPO method to optimize the GQS model, aiming to mitigate the accuracy loss caused by group quantization and pruning by adjusting the weight parameters within each block. Traditional QAT methods typically optimize the entire network’s weights in an end-to-end manner, as described by Equation (2) and (3). Similarly, most pruning methods adopt an end-to-end strategy to globally update unpruned parameters. However, these methods often require substantial computing resources and large-scale training datasets. For instance, pruning the LLaMA-65B model using LLM-Pruner and LoRAPruner requires 154GB and 72GB of memory, respectivelyZhang et al. (2024). To enhance optimization efficiency, BQPO employs a block-wise optimization strategy. Previous studies, such as OmniQuant and AffineQuant, have demonstrated that block-wise optimization can significantly reduce training time and memory usage. However, unlike OmniQuant and AffineQuant, which primarily focus on optimizing quantization parameters such as inter-channel smoothing factors and weight clipping thresholds, GQSA experiences more significant performance degradation due to its high sparsity in structured pruning and low-precision quantization. Therefore, BQPO focuses on optimizing the unpruned weights to restore model performance under extreme compression conditions. Due to the block-wise optimization strategy, this approach minimizes additional training costs, requiring only 26GB of memory for training.
3.4 E2E-OQP
Compared to BQPO, E2E-OQP not only emphasizes intra-block optimization but also considers the overall error across the entire network, thereby capturing cross-block information. As shown in Figure 3 (b), E2E-OQP differs from traditional QAT methods. We assume that BQPO has already produced a well-optimized model in the first stage; thus, E2E-OQP uses BQPO-optimized weights as initialization. During training, we freeze the main network weights $\tilde{W}$ and optimize only the quantization parameters $s$ and $z$ to further fine-tune model performance. The implementation of E2E-OQP highlights the advantages of the GQSA approach. When fine-tuning the quantization parameters, using the BSR format, we quantize the remaining group weights to low bits and freeze them, while the pruned groups are no longer retained. This strategy allows us to effectively fine-tune the quantization parameters and restore the performance of the GQSA model without introducing sparse masks. Overall, E2E-OQP significantly reduces memory usage by focusing exclusively on optimizing the quantization parameters of the remaining group and maintaining the main network at 4-bit quantization. This approach requires only 64GB of memory, which is substantially less than the memory demanded by LoRAPrune.
Figure 4: A simplified view of GQSA’s operator calculation flow. G represents sparse and quantized group size.
3.5 Custom Software Engine
GPU has many processing elements called Streaming Multiprocessors (SMs) and uses a large number of threads to perform computing tasks in parallel. Threads are structured into thread blocks (CTAs), which become the smallest scheduling execution unit on SMs. Therefore, the computation target is decomposed and mapped to each thread block, called CTA, to achieve parallel computing. As shown in Figure 4, for a GEMV task of shape 1×N×K, each thread block is responsible for computing a 1×BN output tile, which is decomposed into $\frac{K}{BK}$ sub-GEMV tasks of shape 1×BN×BK. In offline pre-processing, quantized weights are grouped by size G and saved as gguf format along with scaling factors and zero points. This means that each sub-GEMV task computes $\frac{BK}{G}$ $BN$ non-sparse groups held by one or more output channels. It should be noted that the logical addresses between non-sparse groups are not necessarily consecutive, so the corresponding activation group needs to be accessed according to the real group index of each group. 1 The thread-block issues asynchronous copy instructions to fetch small chunks of input data (tiles) from global memory to shared memory. 2 As soon as a tile arrives in shared memory, it is further sliced into smaller chunks (fragments) and copied into registers. 3 Once all necessary components are in the registers, the quantized matrix undergoes dequantization. 4 The dequantized matrix and inputs are then processed by TensorCores (MMA) or CudaCores (FMA) instructions. 5 Finally, the accumulated results are written back from the registers to the outputs in global memory.
Figure 5: Workload balancing through parallel task partitioning.
LLaMA-7B LLaMA-13B LLaMA-2-7B LLaMA-2-13B LLaMA-3-8B LLaMA-3.1-8B
Setting Method WikiText2 C4 WikiText2 C4 WikiText2 C4 WikiText2 C4 WikiText2 C4 WikiText2 C4
GPTQ 44.01 27.71 15.60 15.29 36.77 33.70 28.14 20.97 210 4.1e4 250 80.3
QUIP 29.74 33.74 12.48 21.94 39.73 31.94 13.48 16.16 84.97 1.3e2 - -
PB-LLM 24.61 49.73 17.73 26.93 25.37 29.84 49.81 19.82 44.12 79.2 - -
Omniquant 15.47 24.89 13.21 18.31 37.37 90.64 17.21 26.76 2.1e4 6.0e4 7.3e3 1.3e4
LeanQuant 15.65 17.62 9.64 10.93 16.98 17.89 10.32 11.73 41.78 36.50 - -
W2 SliM-LLM 14.58 32.91 8.87 13.85 16.01 16.00 9.41 9.41 39.66 1.1e2 - -
2:4 SparseGPT 11.20 13.59 9.14 11.34 10.95 13.56 8.32 11.30 16.56 22.99 16.62 23.22
2:4 Wanda 11.53 14.41 9.58 12.07 11.02 15.07 8.27 12.12 25.27 36.40 23.93 36.24
w4s20% 6.58 8.30 5.75 7.57 6.57 8.32 5.86 7.51 8.43 12.54 8.40 12.37
w4s30% 7.91 9.74 6.72 8.32 7.56 9.49 6.87 8.49 9.79 14.58 9.69 14.32
w4s40% 9.10 11.24 7.70 9.57 8.43 11.31 7.13 9.53 11.80 17.61 11.56 17.32
w4s50% GQSA 11.33 12.03 9.21 10.85 10.64 12.82 7.80 10.93 13.81 20.85 13.56 20.43
Table 1: Wikitext2 and C4 perplexity (↓) for LLaMA-1, LLaMA-2, LLaMA-3 and LLaMA-3.1 models, with a context length of 2048.
Furthermore, to enhance the efficiency of sparse computing, we introduced Stream-KOsama et al. (2023). As show in Figure 5, the classic Slice-KGuo et al. (2024) assigns output tiles independently to thread blocks. Each thread block processes one or more rows of the left operand and one or more columns of the right operand to compute the corresponding output tile by slicing along the internal K dimensions. However, when the weight matrix exhibits high sparsity, the uneven distribution of workloads can result in the "straggler" problem, where small workloads cause inefficiencies. Stream-K addresses this issue by decomposing the workload at a finer granularity, allowing multiple thread blocks to collaborate in computing a single output tile.
4 Experiments
In this section, we provide a comprehensive evaluation of GQSA, focusing on both model performance and inference speed, and demonstrate its advantages over existing model compression techniques through extensive experiments.
4.1 Experimental Settings
Models and Tasks. We selected the LLaMA Touvron et al. (2023a), LLaMA-2Touvron et al. (2023b), LLaMA-3, LLaMA-3.1Dubey et al. (2024) and OPTZhang et al. (2022) models to benchmark our method. Following previous studies, we evaluated the model’s language modeling capability on the WikiText2Merity et al. (2016) and C4Raffel et al. (2020) datasets. To assess performance on zero-shot tasks, we selected several mainstream benchmarks, including PIQABisk et al. (2020), ARCClark et al. (2018), HellaSwagZellers et al. (2019), and WinograndeSakaguchi et al. (2021), and conducted evaluations using lm-eval [^1] https://github.com/EleutherAI/lm-evaluation-harness.
Baselines. We conducted a comprehensive comparison of our method with several recently published techniques in both structured and semi-structured pruning. Given that our implementation achieved INT4 along with 50% structured pruning, we also compared our approach with pure INT2 quantization. For structured pruning, we compared our results with LLMPrunerMa et al. (2023), SliceGPTAshkboos et al. (2024) and ShortGPTMen et al. (2024). For semi-structured pruning, we utilized SparseGPTFrantar and Alistarh (2023) and WandaSun et al. (2023) for comparison. Additionally, we selected OmniQuantShao et al. (2023), QUIPChee et al. (2024), PB-LLMShang et al. (2023), GPTQFrantar et al. (2022), LeanQuantZhang and Shrivastava (2024), and SliM-LLMHuang et al. (2024) as benchmarks for W2 quantization.
Implementation Details. To evaluate the performance of GQSA across various configurations, we implemented sparsity levels of 20%, 30%, 40%, and 50%, using 4-bit weight-only per-group quantization. To strike a balance between model performance and inference speed, a group size of 16 was selected as the optimal configuration. The AdamW optimizerLoshchilov (2017) with a learning rate of 1e-5 was employed to optimize both BQPO and E2E-OQP. The optimization data was randomly sampled from the WikiText2 and C4 datasets, consisting of 4,096 samples, each containing 2,048 tokens. BQPO was trained for 5 epochs, while E2E-OQP was trained for 2 epochs.
4.2 Evaluation on Language Generation Tasks
To assess the performance of GQSA under extreme compression conditions, we first compared its perplexity against baseline method. As shown in Table 1, GQSA surpasses the performance of current state-of-the-art weight-only per-group quantization methods, including GPTQ, QUIP, OmniQuant, LeanQuant, under a 50% structured pruning combined with INT4 quantization. It also surpasses mixed-precision quantization models like PB-LLM and SliM-LLM. Furthermore, GQSA achieves comparable results to 2:4 semi-structured pruning while delivering substantial improvements in compression ratio and speedup. Similar results are presented in Table 9 for OPT models (ranging from 1.3B to 13B parameters), where GQSA consistently matches or surpasses baseline methods, even under more stringent compression settings. Furthermore, we observe that existing model compression methods often experience significant performance degradation on the latest large language models (e.g., LLaMA-3 and LLaMA-3.1). In contrast, GQSA demonstrates robust performance even in scenarios where other methods encounter substantial performance degradation.
Model Setting Method PIQA ARC-C ARC-E Hellaswag Winogrande
ShortGPT 60.1 31.0 41.7 44.0 60.8
SliceGPT 67.5 34.5 55.6 55.1 62.9
25% LLM-Pruner 75.7 37.2 62.0 60.1 62.2
W4S30% GQSA 74.32 34.98 66.04 64.40 65.98
ShortGPT 50.7 27.7 25.6 30.1 50.3
SliceGPT 58.5 27.3 43.5 43.6 57.9
40% LLM-Pruner 70.7 31.3 50.7 53.5 56.1
LLaMA-2-7B W4S40% GQSA 71.27 30.72 61.32 58.48 61.48
ShortGPT 73.1 41.9 60.1 60.6 70.5
SliceGPT 69.6 40.2 61.5 59.4 67.0
25% LLM-Pruner 79.4 43.5 67.8 65.4 63.5
W4S30% GQSA 75.68 39.85 71.55 70.45 66.54
ShortGPT 62.4 32.2 44.8 47.8 62.8
SliceGPT 59.9 29.2 44.1 49.6 61.6
40% LLM-Pruner 75.3 35.4 56.3 60.2 57.8
LLaMA-2-13B W4S40% GQSA 75.30 35.82 66.50 65.40 65.98
Table 2: Zero-shot performance between LLaMA-2-7B and LLaMA-2-13B models under 25% and 40% structured pruning, GQSA with 30% and 40% structured pruning along with INT4 quantization.
4.3 Evaluation on Zero-Shot Tasks
To further validate our model, we conducted a detailed comparison of its zero-shot accuracy against baseline methods. Given the limited data availability from these baselines methods, we selected LLaMA-2-7B and LLaMA-2-13B for the analysis. Table 2 compares GQSA with structured pruning, where GQSA achieved substantial performance gains at equivalent or higher pruning rates, with these benefits becoming more pronounced at higher pruning levels. Table 3 compares GQSA with semi-structured pruning and W2 weight-only per-group quantization. Compared to W2 per-group quantization, GQSA consistently delivered superior performance improvements at the same compression ratio. Under the conditions of 50% structured pruning with INT4 quantization, GQSA outperformed OmniQuant W2 per-group quantization, yielding average accuracy gains of 5.4% for LLaMA-2-7B and 5.7% for LLaMA-2-13B. Given that GQSA operates in a more challenging compression setting than semi-structured pruning, we compare GQSA W4 40% with semi-structured pruning. Experimental results reveal that GQSA achieves superior performance even with a compression rate 3$\times$ higher than that of 2:4 pruning. Furthermore, GQSA demonstrates significant advantages in both speed and accuracy compared to 2:4 pruning. Considering its compression efficiency and flexibility, GQSA emerges as the clear superior choice.
Model Setting Method PIQA ARC-C ARC-E Hellaswag Winogrande
OmniQuant 64.52 26.10 44.94 49.27 54.53
W2 LeanQuant 65.4 24.7 44.2 - 57.4
W4S50% GQSA 68.01 29.01 58.33 52.72 58.41
SparseGPT 70.13 29.35 61.14 56.89 63.14
2:4 Wanda 70.12 30.55 61.32 55.34 62.83
LLaMA-2-7B W4S40% GQSA 71.27 30.72 61.32 58.48 61.48
OmniQuant 68.06 30.03 57.07 56.56 52.95
W2 LeanQuant 70.6 28.2 56.7 - 60.7
W4S50% GQSA 72.47 33.28 63.01 62.11 62.28
SparseGPT 72.74 32.59 66.04 62.78 66.54
2:4 Wanda 73.72 34.39 66.33 63.12 66.93
LLaMA-2-13B W4S40% GQSA 75.30 35.82 66.50 65.40 65.98
Table 3: Zero-shot performance between LLaMA-2-7B and LLaMA-2-13B under W2 quantization method, 50% semi-structured pruning, and GQSA with 40% and 50% structured pruning along with INT4 quantization.
Figure 6: Comparison of GEMV acceleration of our GQSKernel on RTX 4080.
4.4 Inference Engine Evaluation
Kernel Benchmark. We compared GQSKernel with the 2:4 sparse kernel on a (1, 4096) × (4096, 4096) dimension. Due to the flexibility of GQSKernel, it can accommodate varying group sparsity sizes. The experimental results, presented in Figure 6, show that as sparsity increases, the GEMV computation speed improves. Moreover, GQSKernel consistently outperforms the 2:4 sparse mode across all group granularity settings. At 50% sparsity, GQSA achieves a 3$\times$ inference speedup compared to the 2:4 sparse mode.
Figure 7: Inference latency (top) and memory usage (bottom) on an NVIDIA A800-40GB GPU with a fixed input length of 15. W8 results are provided in the Appendix Table 10.
End-to-end throughput. The acceleration of quantization primarily results from memory access savings, whereas sparsity acceleration arises from both memory access and computational savings. We integrated the GQSKernel into FastTransformer and compared it with the FP16 implementation. The experimental results, as shown in Figure  7, indicate that GQSA achieves a 4× reduction in inference latency on the LLaMA-7B model under the GQSA W4S50% setting with a 1024 output length. Additionally, as presented in Appendix Table 6, GQSA further enhances the acceleration potential of the compressed model compared to separate quantization or sparsity methods by simultaneously reducing redundancy in both dimensions of LLMs. For instance, the inference speeds of LLaMA-7B for S50%, W2, and W4S50% are 878.90 ms, 475.55 ms, and 377.98 ms, respectively. Overall, GQSA demonstrates the most significant performance improvement.
SeqLen Method Latency (ms)
W4A16 642.24
W4 2:4 Pruning 513.79
128 GQSA W4 S50% 377.98
W4A16 1312.91
W4 2:4 Pruning 1112.96
256 GQSA W4 S50% 699.26
W4A16 2707.26
W4 2:4 Pruning 1966.45
512 GQSA W4 S50% 1433.43
W4A16 5786.8
W4 2:4 Pruning 4118.36
1024 GQSA W4 S50% 3110.54
Table 4: The inference latency and memory usage of GQSA and 2:4 pruning are compared on an NVIDIA A800-40GB GPU with a fixed input length of 15.
Additionally, we compared GQSA’s performance with that of state-of-the-art sparse schemes, such as SparseGPT and Wanda’s 2:4 sparse scheme. The experimental results, presented in Table 4, demonstrate that GQSA outperforms these methods in terms of inference latency and accuracy.
4.5 Ablation Experiments
We investigated the impact of group size and sparsity on the performance of the GQSA model. As shown in Figure 8 (left), GQSA demonstrates robust performance at sparsity levels of 50% or lower. When sparsity exceeds 60%, a noticeable performance degradation occurs. However, even at an extreme sparsity level of 80%, GQSA achieves a perplexity below 30, avoiding performance collapse. Figure 8 (right) illustrates the relationship between group size and model performance. Overall, model performance exhibits a clear correlation with group size. Based on performance considerations, we selected 16 as the default group size for the model.
Figure 8: The ablation studies on the LLaMA-7B model to evaluate the impact of different structured pruning group sizes (right) and sparsity levels (left) on model performance.
5 Conclusion
We propose GQSA, an efficient sparse acceleration method for the decoding process, compatible with weight-only per-group quantization. Through a comprehensive analysis of LLMs weights, we investigated group sparse modes beyond the 2:4 sparsity mode. To enhance model performance, we implemented a two-stage sparse optimization strategy, comprising BQPO and E2E-OQP. Based on the BSR format, we then developed an efficient sparse inference engine to fully leverage the synergistic benefits of quantization and sparsity. Extensive experimental results demonstrate that GQSA effectively integrates at both the algorithmic and system levels, offering a superior accuracy-speed trade-off compared to traditional 2:4 sparsity and quantization approaches.
Limitations
The proposed GQSA extends beyond the 2:4 sparsity pattern to explore group sparsity patterns, enabling efficient compatibility with weight-only per-group quantization. By combining algorithm-level optimizations with a customized inference engine, our approach achieves an improved balance between accuracy and inference speed. However, the current method does not address activation quantization, and due to resource limitations, it has not yet been applied to large language models (LLMs) exceeding 100 billion parameters. These limitations present promising directions for future research, and we are optimistic that they will be addressed in subsequent work.
Ethics Statement
This paper introduces a method to tackle the challenges of compressing large language models (LLMs), with the goal of facilitating their wider application and adoption. In the context of current research, ethical considerations surrounding LLMs have received substantial attention. Our findings indicate that the proposed method does not exacerbate existing biases or compromise ethical standards.
Appendix
Appendix A The effects of BQPO and E2E-OQP on the model’s performance
Table 5 presents the impact of BQPO and E2E-OQP on model performance. BQPO optimizes weights in a block-wise manner, effectively preserving the performance of the GQS model. Finally, E2E-OQP, which accounts for cross-layer errors, yields the best model performance.
Method LLaMA-13B LLaMA-2-13B
WikiText2 C4 WikiText2 C4
BQPO 12.90 13.39 10.55 13.56
BQPO+E2E-OQP 9.21 10.85 7.80 10.93
Table 5: The effectiveness of BQPO and E2E-OQP methods for compressing LLaMA-13B and LLaMA-2-13B models.
Appendix B A comparison of the effects of pruning and quantization on inference performance
Setting WikiText2 C4 Inference speed (ms)
0% 5.47 6.97 1490.50
S20% 7.67 9.10 1370.35
S30% 9.34 11.27 1181.25
S40% 10.84 16.38 1035.15
S50% 14.56 21.09 878.90
S60% 25.76 37.49 671.98
W8 5.50 7.01 868.35
W4 5.72 7.25 642.24
W2 36.43 40.34 475.55
W4S50% 10.64 12.82 377.98
Table 6: Performance comparison of GQSA with naive pruning and naive quantization in the extreme compression setting on LLaMA-2-7B.
As demonstrated in Table 6, we will highlight the comprehensive performance advantages of GQSA over single pruning and quantization methods from two perspectives.
From the Perspective of Algorithm Accuracy: Both quantization and sparsity, when applied individually, can lead to significant accuracy degradation under extreme compression settings. For instance, the PPL test results under S60% and W2 configurations demonstrate considerable performance loss. However, combining these two strategies allows for higher compression rates while better preserving model performance compared to using either strategy alone. As an example, using the LLaMA-2-7B WikiText2 benchmark, the results for W2, S60%, and W4S50% are 36.44, 25.76, and 10.64, respectively.
From the Perspective of Inference Speed: The acceleration benefit of quantization primarily arises from reduced memory access, while the acceleration benefit of sparsity stems from both memory and computational savings. For pure quantization or pure sparsity, the acceleration benefit diminishes as the compression rate increases. GQSA, however, enhances the upper limit of the acceleration benefit by simultaneously reducing redundancy in both dimensions (quantization and sparsity). For example, in the case of LLaMA-2-7B, the inference speeds for S60%, W2, and W4S50% are 671.98, 475.55, and 377.98, respectively.
Appendix C Combining the advantages of structured pruning and group quantization
As show in Table 7 the acceleration benefits of quantization primarily stem from reduced memory access, while sparsity accelerates inference by saving both memory and computation (as sparse groups do not need to be stored, read, or computed). When applying only the quantization strategy, the LLM’s acceleration benefit does not increase exponentially as the bit-width of $W$ decreases. Instead, it faces diminishing returns, as the performance bottleneck shifts from memory access to computation as the quantization bit-width is reduced. In contrast, GQSA can further accelerate deep quantization models by skipping redundant calculations, thereby pushing the upper limit of acceleration benefits. For example, in the case of LLaMA-2-7B, the measured inference speed is 20% faster with W4S50 than with W2.
Model Setting Inference speed (ms)
LLaMA-2-7B W4 642.24
W2 475.55
W4S50% 377.98
Table 7: Comparison of inference speed between GQSA and single quantization.
Appendix D Inference throughput of GQSA
As show in Table 8, we evaluated the throughput of the GQSA model based on FastTransformer on an Nvidia A100 80 GB GPU. The results demonstrate that, compared to the pure W8 and W4 configurations, GQSA’s W8S50% and W4S50% configurations achieved a 60% improvement in throughput.
Setting LLaMA-7B LLaMA-13B
FP 92.69 50.68
W8 156.40 95.78
W8S50 263.64 158.99
W4 202.81 137.92
W4S50 343.43 228.95
Table 8: Inference throughput (tokens per second) of GQSA on the NVIDIA A100 80GB.
Appendix E Differences from Sparse Methods in Traditional CNNs
Although previous work, such as PatDNN, introduced semi-structured sparsity in CNN networks, we believe our work contributes to the field in two core aspects: First, we have significantly advanced the engineering implementation of semi-structured sparsity. Notably, we introduced the "task-centric" parallel strategy, replacing the widely-used "data-centric" parallel approach in the industry. This shift effectively addresses the issue of unbalanced load across computing units, resulting in a substantial speedup of 1.3$\times$ to 1.5$\times$ for individual operators, thus achieving a new state-of-the-art in engineering performance. Second, while the GEMM operator in traditional CNN networks typically adopts the "N×$\times$1" sparse mode, we propose the "1$\times$×N" sparse mode tailored to the characteristics of LLM models. This innovation better preserves outliers within the channel and is fundamentally different from the traditional "N×$\times$1" mode in terms of engineering implementation.
We believe innovation is not solely about proposing "new concepts" or "new strategies" but also about selecting the most appropriate approaches to address real technical challenges and pushing the performance boundaries. Currently, the LLM field faces significant inference cost challenges, and relying exclusively on quantization techniques has nearly reached its performance optimization limits. Our work contributes to further enhancing performance based on quantization models and has led to a SOTA breakthrough in semi-structured sparsity technology within the LLM field. The pursuit of higher performance limits and greater industrial applicability reflects a key aspect of innovation.
Appendix F The advantages of group quantization compared to standard per-layer or per-label quantization methods.
From the Perspective of Quantization Accuracy: The primary challenge in quantization LLMs arises from the imbalanced numerical distribution (both between and within channels) and the prevalence of outliers in both weights and activations. Standard per-layer and per-token (or per-channel) quantization methods assume that the entire tensor or the neurons in each channel are identically distributed. This coarser quantization granularity is insufficient to address the issues of uneven distribution and outlier retention. Group quantization, however, further partitions the channels and quantizes the model weights at a finer granularity, effectively mitigating the problem of imbalanced numerical distribution and improving outlier handling, thereby reducing the accuracy loss typically associated with quantization.
From the Perspective of Quantization Speed: The finer quantization granularity of the per-group approach necessitates additional scaling factors during computation. However, since LLM tasks are memory-intensive rather than computation-bound, this increased granularity does not significantly impact memory access complexity compared to 2:4 sparsity. As a result, the inference speed is not adversely affected. For instance, the widely used reasoning engine, llama.cpp, employs group quantization for model inference.
Appendix G Results of GQSA on the OPT model
To verify the generalization ability of GQSA on different model families, we conduct experiments on OPT models (ranging from 1.3B to 13B parameters). Table  9 shows similar results to the LLaMA model family, where GQSA consistently matches or outperforms the baseline methods even under stricter compression settings.
OPT-1.3B OPT-2.7B OPT-6.7B OPT-13B
Setting Method WikiText2 C4 WikiText2 C4 WikiText2 C4 WikiText2 C4
GPTQ 130.88 60.88 61.59 33.83 20.18 18.55 21.36 16.34
QUIP 41.64 - 28.98 - 18.57 - 16.02 -
PB-LLM 45.92 - 39.71 - 20.37 - 19.11 -
OmniQuant 23.95 27.33 18.13 21.11 14.43 16.67 12.94 14.92
W2 SliM-LLM 24.57 - 17.98 - 14.22 - 12.16 -
SparseGPT 24.54 26.55 17.82 19.45 14.23 16.56 12.94 14.88
2:4 Wanda 28.27 28.54 21.17 22.84 15.90 18.99 15.55 16.18
W4S20% 14.49 16.60 12.03 14.54 10.21 12.71 9.93 12.16
W4S30% 16.06 18.44 13.23 15.95 10.94 13.64 10.37 12.85
W4S40% 18.82 21.54 15.39 18.25 12.15 15.12 11.29 13.97
W4S50% GQSA 21.32 24.90 17.52 20.81 13.44 16.94 12.16 15.57
Table 9: Wikitext2 and C4 perplexity (↓) for OPT models, with a context length of 2048.
Appendix H GQSA inference latency and memory consumption
Due to space constraints, detailed inference latency and model memory consumption are provided in Appendix Table 10. Overall, GQSA demonstrates exceptional performance across various settings.
LLaMA-7B
128 256 512 1024
sequence length Latency (ms) Memory (GB) Latency (ms) Memory (GB) Latency (ms) Memory (GB) Latency (ms) Memory (GB)
fp16 1490.5 13.47 3005.95 13.534 6090.97 13.662 12561.82 13.918
w8a16 868.35 7.394 1755.62 7.458 3594.95 7.586 7559.22 7.842
w8a16+sp0.3 688.89 6.296 1261.05 6.361 3005.02 6.489 5814.62 6.745
w8a16+sp0.4 603.23 5.669 1103.08 5.733 2593.76 5.861 5039.33 6.117
w8a16+sp0.5 512.71 5.042 996.59 5.106 2019.1 5.234 4329.32 5.492
w4a16 642.24 4.258 1312.91 4.322 2707.26 4.45 5786.8 4.706
w4a16+g16+sp0.3 518.99 4.101 1041.18 4.165 2113.56 4.293 4437.48 4.549
w4a16+g16+sp0.4 432.05 3.788 855.46 3.852 1828.48 3.977 3772.63 4.233
w4a16+g16+sp0.5 377.98 3.474 699.26 3.528 1433.43 3.653 3110.54 3.909
LLaMA-13B
128 256 512 1024
sequence length Latency (ms) Memory (GB) Latency (ms) Memory (GB) Latency (ms) Memory (GB) Latency (ms) Memory (GB)
fp16 2726.66 25.61 5481.96 25.696 11071.81 25.92 22559.77 26.304
w8a16 1439.66 13.524 2900.46 13.62 5922.28 13.844 12257.27 14.228
w8a16+sp0.3 1164.239 11.396 2114.165 11.492 4976.471 11.716 9501.55 12.105
w8a16+sp0.4 1024.199 10.182 1843.61 10.278 4272.72 10.502 8343.77 10.886
w8a16+sp0.5 869.486 8.964 1715.976 9.061 3345.762 9.285 7044.25 9.669
w4a16 999.1 7.444 2020.99 7.54 4155.94 7.764 8750.98 8.148
w4a16+g16+sp0.3 801.203 7.141 1475.175 7.237 3563.465 7.461 6972.12 7.845
w4a16+g16+sp0.4 702.602 6.532 1303.865 6.628 3087.623 6.852 6081.292 7.236
w4a16+g16+sp0.5 603.515 5.924 1104.366 6.02 2374.286 6.244 5099.638 6.628
LLaMA-30B (TP=2)
128 256 512 1024
sequence length Latency (ms) Memory (GB) Latency (ms) Memory (GB) Latency (ms) Memory (GB) Latency (ms) Memory (GB)
fp16 3759.08 65.534 7540.17 65.726 15241.36 66.11 31073.23 66.878
w8a16 3032.64 32.418 6111.43 32.642 12371.66 33.026 25477.58 33.794
w8a16+g16+sp0.3 2412.6 27.084 4861.575 27.308 10343.645 27.692 19826.459 28.46
w8a16+g16+sp0.4 2132.65 24.036 3840.98 24.261 8925.685 24.645 17343.09 25.413
w8a16+g16+sp0.5 1797.27 20.988 3472.16 21.212 6950 21.596 14591.638 22.364
w4a16 1938.2 17.178 3924.2 17.402 8011.57 17.786 16680.64 18.554
w4a16+g16+sp0.3 1541.925 16.416 2515.512 16.641 6800.993 17.025 13290.836 17.793
w4a16+g16+sp0.4 1341.315 14.892 2229.65 15.116 5890.861 15.501 11591.38 16.269
w4a16+g16+sp0.5 1122.292 13.368 2180.11 13.592 4526.311 13.816 9720.279 14.584
Table 10: Inference latency and memory usage of the FastTransformer implementation on NVIDIA A800-40GB GPU with a fixed input sequence length of 15, output sequence lengths of 128, 256, 512 and 1024.

### v4

@@
 Model compression has emerged as a mainstream solution to reduce memory usage and computational overhead.
-This paper presents Group Quantization and Sparse Acceleration (GQSA), a novel compression technique tailored for LLMs.
+This paper proposes GQSA, a novel model compression framework specifically designed for LLMs.
 Traditional methods typically focus exclusively on either quantization or sparsification, but relying on a single strategy often results in significant performance loss at high compression rates.
@@
 Freezing the backbone network can not only reduce memory usage but also improve training efficiency.
-Extensive experiments demonstrate that GQSA achieves significant advantages in both model accuracy and inference speed, especially when applied to newly released advanced models such as LLaMA-3 and LLaMA-3.1 model family.
+Extensive experiments demonstrate that GQSA achieves significant advantages in both model accuracy and inference speed, especially when applied to newly released advanced models such as LLaMA-3 and LLaMA-3.1 model family and Qwen2.5 models.
 In summary, our contributions are as follows.
-  • We propose a sparse scheme seamlessly compatible with widely used weight-only per-group quantization, effectively accelerating GEMV operations and reducing memory usage.
+  • We propose a sparse scheme seamlessly compatible with widely used weight-only and weight-activation quantization, effectively accelerating GEMV operations and reducing memory usage.
   • We introduce a task-centric parallel implementation, addressing the workload balancing issue in sparse acceleration.
@@
 Figure 1: The distribution of the top 1% significant weights in the Hessian matrix, derived from the $k_{proj}$ and $q_{proj}$ distributions in the LLaMA-7B model.
+Figure 2: Overview of GQSA.
+(a) We propose a group quantization and sparse LLMs, where linear layers are replaced by GQS layers.
+(b) We use the two-stage optimization method BQPO and E2E-OQP to recover the performance of the extremely compressed model.
 Advantages of GQSA.
@@
 3 GQSA
+In this section, we provide a detailed exposition of GQSA.
+Section 3.1 begins with an examination of weight quantization and salient weight selection principles.
+Building upon these foundations, Section 3.2 introduces the innovative GQS Layer, designed to maximize the compression advantages from both quantization and pruning.
+The subsequent sections ( 3.3 and  3.4) detail our two-stage optimization algorithm, which delivers exceptional model performance.
+Concluding the section,  3.5 proposes a novel task-centric parallel strategy for efficient inference acceleration.
 3.1 Preliminary
@@
 Group pruning is then applied based on the average saliency metrics within each group, resulting in the formation of the GQS layer.
-Figure 3: Overview of GQSA.
-(a) We propose a group quantization and sparse LLMs, where linear layers are replaced by GQS layers.
-(b) We use the two-stage optimization method BQPO and E2E-OQP to recover the performance of the extremely compressed model.
 3.2 GQS Layer
@@
 -1, 7, 14, 0, 0, 3, 6, 15}
-where rowIndex[i] represent the offset of each row i, where i belongs to the range $[0,\text{rows}]$.
-The difference rowIndex[r+1] - rowIndex[i] indicates the number of non-zero groups in the i-th row.
-Additionally, rowIndex[rows] represents the total number of non-zero groups.
-The array groups[i] stores the indices of the non-zero groups; for instance, if groups[1] = 0, it means that the second group is located in the 0th column (in terms of group units).
-Finally, values stores the values of the non-zero groups for each row.
+[…]
 3.3 BQPO
-In the first stage (shown in Figure 3(b)), we apply the BQPO method to optimize the GQS model, aiming to mitigate the accuracy loss caused by group quantization and pruning by adjusting the weight parameters within each block.
-Traditional QAT methods typically optimize the entire network’s weights in an end-to-end manner, as described by Equation (2) and (3).
-Similarly, most pruning methods adopt an end-to-end strategy to globally update unpruned parameters.
-However, these methods often require substantial computing resources and large-scale training datasets.
-For instance, pruning the LLaMA-65B model using LLM-Pruner and LoRAPruner requires 154GB and 72GB of memory, respectivelyZhang et al.
-(2024).
-To enhance optimization efficiency, BQPO employs a block-wise optimization strategy.
-Previous studies, such as OmniQuant and AffineQuant, have demonstrated that block-wise optimization can significantly reduce training time and memory usage.
-However, unlike OmniQuant and AffineQuant, which primarily focus on optimizing quantization parameters such as inter-channel smoothing factors and weight clipping thresholds, GQSA experiences more significant performance degradation due to its high sparsity in structured pruning and low-precision quantization.
-Therefore, BQPO focuses on optimizing the unpruned weights to restore model performance under extreme compression conditions.
-Due to the block-wise optimization strategy, this approach minimizes additional training costs, requiring only 26GB of memory for training.
+In the first stage (Figure 2(b)), we apply the BQPO method to optimize the GQS model, aiming to mitigate the accuracy degradation caused by group quantization and pruning.
+This is achieved by adjusting the weight parameters within each block.
+Traditional QAT methods typically optimize the entire network’s weights in an end-to-end fashion, as illustrated in Equations (2) and (3).
+Similarly, most pruning approaches adopt a global end-to-end strategy to update the remaining unpruned parameters.
+However, such methods often demand substantial computational resources and large-scale training datasets.
+To improve optimization efficiency, BQPO adopts a block-wise optimization strategy.
+Prior studies, such as OmniQuant and AffineQuant, have shown that block-wise optimization can significantly reduce both training time and memory consumption.
+Unlike OmniQuant and AffineQuant, which primarily optimize quantization parameters (inter-channel smoothing factors and weight clipping thresholds), GQSA suffers from more severe performance degradation due to its combination of high structured sparsity and low-precision quantization.
+As a result, BQPO focuses on optimizing the remaining weights to recover performance under extreme compression settings.
+This block-wise approach enables significant performance restoration with minimal additional training cost compared to global optimization techniques.
+Figure 4: A simplified view of GQSA’s operator calculation flow. G represents sparse and quantized group size.
 3.4 E2E-OQP
-Compared to BQPO, E2E-OQP not only emphasizes intra-block optimization but also considers the overall error across the entire network, thereby capturing cross-block information.
-As shown in Figure 3 (b), E2E-OQP differs from traditional QAT methods.
-We assume that BQPO has already produced a well-optimized model in the first stage; thus, E2E-OQP uses BQPO-optimized weights as initialization.
-During training, we freeze the main network weights $\tilde{W}$ and optimize only the quantization parameters $s$ and $z$ to further fine-tune model performance.
-The implementation of E2E-OQP highlights the advantages of the GQSA approach.
-When fine-tuning the quantization parameters, using the BSR format, we quantize the remaining group weights to low bits and freeze them, while the pruned groups are no longer retained.
-This strategy allows us to effectively fine-tune the quantization parameters and restore the performance of the GQSA model without introducing sparse masks.
-Overall, E2E-OQP significantly reduces memory usage by focusing exclusively on optimizing the quantization parameters of the remaining group and maintaining the main network at 4-bit quantization.
-This approach requires only 64GB of memory, which is substantially less than the memory demanded by LoRAPrune.
-Figure 4: A simplified view of GQSA’s operator calculation flow. G represents sparse and quantized group size.
+Compared to BQPO, E2E-OQP not only performs intra-block optimization but also accounts for the overall error across the entire network, thereby capturing cross-block dependencies.
+As illustrated in Figure 2(b), E2E-OQP differs from conventional quantization-aware training (QAT) methods.
+Assuming that BQPO has already yielded a well-optimized model in the first stage, E2E-OQP initializes training using the BQPO-optimized weights.
+During this phase, we freeze the primary network weights $\tilde{W}$ and optimize only the quantization parameters $s$ and $z$ to further refine model performance.
+The design of E2E-OQP underscores the advantages of the GQSA framework.
+Specifically, during fine-tuning, we employ the block-sparse row (BSR) format: the remaining group weights are quantized to low bit-width and frozen, while pruned groups are discarded entirely.
+This strategy enables effective fine-tuning of the quantization parameters without requiring sparse masks, thereby restoring the performance of the GQSA model under extreme compression.
+Overall, E2E-OQP achieves substantial memory savings by focusing solely on the quantization parameters of the remaining groups while maintaining 4-bit quantization across the main network.
+A detailed comparison of the resource consumption of BQPO and E2E-OQP is provided in Appendix A, demonstrating the efficiency advantages of the GQSA approach.
+Figure 5: Workload balancing through parallel task partitioning.
 3.5 Custom Software Engine
@@
 5 Finally, the accumulated results are written back from the registers to the outputs in global memory.
-Figure 5: Workload balancing through parallel task partitioning.
 LLaMA-7B LLaMA-13B LLaMA-2-7B LLaMA-2-13B LLaMA-3-8B LLaMA-3.1-8B
@@
 2:4 Wanda 11.53 14.41 9.58 12.07 11.02 15.07 8.27 12.12 25.27 36.40 23.93 36.24
-w4s20% 6.58 8.30 5.75 7.57 6.57 8.32 5.86 7.51 8.43 12.54 8.40 12.37
-w4s30% 7.91 9.74 6.72 8.32 7.56 9.49 6.87 8.49 9.79 14.58 9.69 14.32
-w4s40% 9.10 11.24 7.70 9.57 8.43 11.31 7.13 9.53 11.80 17.61 11.56 17.32
-w4s50% GQSA 11.33 12.03 9.21 10.85 10.64 12.82 7.80 10.93 13.81 20.85 13.56 20.43
+[…]
+[…]
+[…]
+[…]
 Table 1: Wikitext2 and C4 perplexity (↓) for LLaMA-1, LLaMA-2, LLaMA-3 and LLaMA-3.1 models, with a context length of 2048.
@@
 4 Experiments
-In this section, we provide a comprehensive evaluation of GQSA, focusing on both model performance and inference speed, and demonstrate its advantages over existing model compression techniques through extensive experiments.
 4.1 Experimental Settings
@@
 (2019), and Winogrande Sakaguchi et al.
-(2021), and conducted evaluations using lm-eval [^1] https://github.com/EleutherAI/lm-evaluation-harness.
+(2021), and conducted evaluations using lm-eval.
 Baselines.
@@
 Furthermore, GQSA achieves comparable results to 2:4 semi-structured pruning while delivering substantial improvements in compression ratio and speedup.
-Similar results are presented in Table 9 for OPT models (ranging from 1.3B to 13B parameters), where GQSA consistently matches or surpasses baseline methods, even under more stringent compression settings.
+Similar results are presented in Table 14 for Qwen2.5 and Table 15 for OPT models, where GQSA consistently matches or surpasses baseline methods, even under more stringent compression settings.
 Furthermore, we observe that existing model compression methods often experience significant performance degradation on the latest large language models (e.g., LLaMA-3 and LLaMA-3.1).
@@
 25% LLM-Pruner 75.7 37.2 62.0 60.1 62.2
-W4S30% GQSA 74.32 34.98 66.04 64.40 65.98
+[…]
 ShortGPT 50.7 27.7 25.6 30.1 50.3
@@
 40% LLM-Pruner 70.7 31.3 50.7 53.5 56.1
-LLaMA-2-7B W4S40% GQSA 71.27 30.72 61.32 58.48 61.48
+[…]
 ShortGPT 73.1 41.9 60.1 60.6 70.5
@@
 25% LLM-Pruner 79.4 43.5 67.8 65.4 63.5
-W4S30% GQSA 75.68 39.85 71.55 70.45 66.54
+[…]
 ShortGPT 62.4 32.2 44.8 47.8 62.8
@@
 40% LLM-Pruner 75.3 35.4 56.3 60.2 57.8
-LLaMA-2-13B W4S40% GQSA 75.30 35.82 66.50 65.40 65.98
+[…]
 Table 2: Zero-shot performance between LLaMA-2-7B and LLaMA-2-13B models under 25% and 40% structured pruning, GQSA with 30% and 40% structured pruning along with INT4 quantization.
@@
 W2 LeanQuant 65.4 24.7 44.2 - 57.4
-W4S50% GQSA 68.01 29.01 58.33 52.72 58.41
+[…]
 SparseGPT 70.13 29.35 61.14 56.89 63.14
 2:4 Wanda 70.12 30.55 61.32 55.34 62.83
-LLaMA-2-7B W4S40% GQSA 71.27 30.72 61.32 58.48 61.48
+[…]
 OmniQuant 68.06 30.03 57.07 56.56 52.95
 W2 LeanQuant 70.6 28.2 56.7 - 60.7
-W4S50% GQSA 72.47 33.28 63.01 62.11 62.28
+[…]
 SparseGPT 72.74 32.59 66.04 62.78 66.54
 2:4 Wanda 73.72 34.39 66.33 63.12 66.93
-LLaMA-2-13B W4S40% GQSA 75.30 35.82 66.50 65.40 65.98
+[…]
 Table 3: Zero-shot performance between LLaMA-2-7B and LLaMA-2-13B under W2 quantization method, 50% semi-structured pruning, and GQSA with 40% and 50% structured pruning along with INT4 quantization.
@@
 W4 2:4 Pruning 513.79
-128 GQSA W4 S50% 377.98
+[…]
 W4A16 1312.91
 W4 2:4 Pruning 1112.96
-256 GQSA W4 S50% 699.26
+[…]
 W4A16 2707.26
 W4 2:4 Pruning 1966.45
-512 GQSA W4 S50% 1433.43
+[…]
 W4A16 5786.8
 W4 2:4 Pruning 4118.36
-1024 GQSA W4 S50% 3110.54
+[…]
 Table 4: The inference latency and memory usage of GQSA and 2:4 pruning are compared on an NVIDIA A800-40GB GPU with a fixed input length of 15.
@@
 Appendix
-Appendix A The effects of BQPO and E2E-OQP on the model’s performance
+Appendix A Training Efficiency of GQSA
+Table 5 lists the memory and time required to train the Lllama-2 model using GQSA.
+The results show that GQSA requires only minimal resource overhead, and the 7B model only takes less than 10 hours to train with 9.3GB of memory, which is much less than the 14GB memory requirement to load the FP16 model, which is very efficient.
+It is also significantly better than other pruning methods, such as LLM-Pruner, which requires 18GB and takes less training time.
+LLaMA-2 BQPO E2E-OQP
+Memory (GB) Time (h) Memory (GB) Time (h)
+7B 9.3 5.1 7.6 4.2
+13B 14.3 7.3 11.7 6.4
+Table 5: Detailed training time and training memory for GQSA at different model sizes and quantization bits on a single A100-40GB GPU.
+Appendix B The effects of BQPO and E2E-OQP on the model’s performance
 Table 6 presents the impact of BQPO and E2E-OQP on model performance.
@@
 Table 6: The effectiveness of BQPO and E2E-OQP methods for compressing LLaMA-13B and LLaMA-2-13B models.
-Appendix B A comparison of the effects of pruning and quantization on inference performance
+Appendix C GQSA performance under weight-activation quantization
+Unlike other algorithms that are limited to weight-only quantization, weight-activation quantization or model pruning, GQSA can not only efficiently combine pruning with weight-only quantization, but also support pruning with weight activation quantization.
+Our GPU-friendly grouped semi-structured sparse solution can be seamlessly combined with weight-only quantization or weight+activation quantization.
+On the basis of quantization, we can further improve performance by skipping some operations through sparsity.
+Model Settings WikiText2 C4
+LLaMA-2-7B W4A8S50% 7.84 11.04
+LLaMA-2-13B W4A8S50% 14.09 21.26
+Table 7: Performance comparison of GQSA with weight-activation quantization.
+As show in Table 7, GQSA effectively preserves model accuracy under W4A8S50% quantization for both LLaMA-2-7B and LLaMA-2-13B architectures, maintaining strong performance despite simultaneous weight and activation quantization with 50% sparsity.
+Appendix D Comparison of Quantized & Pruned Works
+A comparison with SparseGPT’s joint sparsification and quantization.
+In SparseGPT’s report, "Joint Sparsification & Quantization" performs worse than "Sparsification-only," so we initially did not include it in our main content.
+However, for completeness, the following Table 8 presents a direct comparison between GQSA and SparseGPT’s "Joint Sparsification & Quantization" on LLaMA-2-13B and LLaMA-3-8B.
+The results demonstrate that GQSA provides a more significant performance advantage.
+Method LlaMA-2-13B LLaMA-3-8B
+WikiText2 C4 WikiText2 C4
+SparseGPT 2:4 8.32 11.30 16.56 22.99
+SparseGPT 2:4+INT4 9.25 12.74 19.43 26.34
+GQSA W4S50% 7.80 10.93 13.84 20.85
+Table 8: Performance comparison of GQSA with SparseGPT.
+Comparison with contemporaneous works.
+To further validate the superiority of GQSA, we conduct comparative evaluations with contemporaneous methods including SliM-LoRA Mozaffari and Dehnavi (2024) and DC-W8A8 Wang et al.
+(2025).
+SliM-LoRA employs 4-bit weight quantization combined with Wanda’s 2:4 pruning but fails to overcome the limitations of semi-structured sparsity.
+Since NVIDIA’s 2:4 Tensor Cores do not support weight-only quantization, the inference acceleration benefits remain limited.
+Additionally, the 2:4 sparse format retains randomly positioned neurons and requires storing an equal amount of metadata to record their locations, preventing effective memory compression.
+SliM-LoRA also introduces the LoRA-Adapter; however, due to the quantization and sparsity of the main network, the LoRA-Adapter cannot be directly integrated and must be stored separately, increasing inference complexity.
+According to the SliM paper, its sparse quantization matrix can even reduce inference speed on the A100 GPU.
+In contrast, GQSA’s sparse quantization achieves a 4.3 $\times$ inference speedup over FP16, highlighting SliM’s shortcomings in both inference acceleration and memory compression.
+DC-W8A8 incorporates sparsity into W8A8 quantization but relies on unstructured sparsity with a sparsity rate of only 20%, offering minimal memory compression benefits.
+As stated in its paper, DC-W8A8 achieves only a 2.2$\times$ compression ratio compared to FP16, whereas GQSA achieves a 4.3$\times$ compression ratio.
+Moreover, GQSA significantly outperforms DC-W8A8 in inference acceleration.
+Method OPT LLaMA-2
+6.7B 13B 7B 13B
+SliM-LoRA 47.08 47.96 54.26 57.85
+DC-W8A8 48.55 - 60.89 -
+GQSA W4S50% 53.26 56.39 59.36 64.96
+Table 9: Performance comparison between GQSA and contemporaneous methods.
+Table 9 presents the comparative evaluation of average accuracy on zero-shot tasks across different methods.
+The experimental results demonstrate that GQSA consistently outperforms both SliM and DC-W8A8 in terms of overall performance.
+Furthermore, GQSA achieves superior acceleration ratios and compression rates, while maintaining competitive accuracy.
+These advantages make GQSA particularly suitable for edge-side inference scenarios, where both computational efficiency and model compactness are critical.
+Appendix E A comparison of the effects of pruning and quantization on inference performance
 Setting WikiText2 C4 Inference speed (ms)
@@
 For example, in the case of LLaMA-2-7B, the inference speeds for S60%, W2, and W4S50% are 671.98, 475.55, and 377.98, respectively.
-Appendix C Combining the advantages of structured pruning and group quantization
+Appendix F Combining the advantages of structured pruning and group quantization
 As show in Table 11 the acceleration benefits of quantization primarily stem from reduced memory access, while sparsity accelerates inference by saving both memory and computation (as sparse groups do not need to be stored, read, or computed).
@@
 Table 11: Comparison of inference speed between GQSA and single quantization.
-Appendix D Inference throughput of GQSA
+Appendix G Comparison of GQSA with Vector Quantization
+Some of the latest low-bit quantization methods, such as AQLM Egiazarian et al.
+(2024) and QuIP# Tseng et al.
+(2024), employ vector quantization (VQ), which differs from uniform quantization techniques like GQSA.
+VQ constructs codebooks by learning the underlying data distribution, enabling better data preservation and potentially higher model performance.
+However, VQ methods rely on pre-trained codebooks (e.g., the E8P codebook used in QuIP# and the multi-codebook scheme in AQLM), which introduce considerable computational overhead during both training and inference.
+This makes them less practical for real-world deployment.
+Method WikiText2 C4 Tokens Per Second
+QuIP# W2 6.06 8.07 71.09
+AQLM W2 5.60 7.47 68.1
+GQSA W4S50% 7.80 10.93 228.95
+Table 12: Comparison between GQSA and Vector Quantization
+In contrast, GQSA combines uniform quantization with high sparsity to enable efficient inference acceleration in practical scenarios.
+As show in Table 12, while it may slightly underperform VQ-based methods like QuIP# and AQLM in terms of accuracy, it significantly outpaces them in inference speed—achieving up to 3.3$\times$ the speed of vector quantization methods under a small accuracy trade-off.
+No single method perfectly balances accuracy and computational efficiency; GQSA prioritizes inference speed, accepting a minor compromise in model accuracy to achieve substantial gains in performance.
+Appendix H Inference throughput of GQSA
 As show in Table 13, we evaluated the throughput of the GQSA model based on FastTransformer on an Nvidia A100 80 GB GPU.
@@
 Table 13: Inference throughput (tokens per second) of GQSA on the NVIDIA A100 80GB.
-Appendix E Differences from Sparse Methods in Traditional CNNs
+Appendix I Differences from Sparse Methods in Traditional CNNs
 Although previous work, such as PatDNN, introduced semi-structured sparsity in CNN networks, we believe our work contributes to the field in two core aspects: First, we have significantly advanced the engineering implementation of semi-structured sparsity.
@@
 The pursuit of higher performance limits and greater industrial applicability reflects a key aspect of innovation.
-Appendix F The advantages of group quantization compared to standard per-layer or per-label quantization methods.
+Appendix J The advantages of group quantization compared to standard per-layer or per-label quantization methods.
 From the Perspective of Quantization Accuracy: The primary challenge in quantization LLMs arises from the imbalanced numerical distribution (both between and within channels) and the prevalence of outliers in both weights and activations.
@@
 For instance, the widely used reasoning engine, llama.cpp, employs group quantization for model inference.
-Appendix G Results of GQSA on the OPT model
+Appendix K Results of GQSA on the Qwen model
+To verify the generalization ability of GQSA on different model families, we conduct experiments on Qwen models (base and instruct model).
+Table  14 shows similar results to the LLaMA model family, where GQSA consistently matches or outperforms the baseline methods even under stricter compression settings.
+Appendix L Results of GQSA on the OPT model
 To verify the generalization ability of GQSA on different model families, we conduct experiments on OPT models (ranging from 1.3B to 13B parameters).
-Table  9 shows similar results to the LLaMA model family, where GQSA consistently matches or outperforms the baseline methods even under stricter compression settings.
+Table  15 shows similar results to the LLaMA and Qwen model family, where GQSA consistently matches or outperforms the baseline methods even under stricter compression settings.
+Qwen2.5-7B Qwen2.5-14B Qwen2.5-7B-Instruct Qwen2.5-14B-Instruct
+Setting Method WikiText2 C4 WikiText2 C4 WikiText2 C4 WikiText2 C4
+GPTQ 29.22 89.12 23.08 55.43 46.03 289.92 48.42 38.37
+W2 OmniQuant 14.49 22.78 11.98 17.81 17.26 26.37 12.95 17.97
+sparsegpt 11.25 17.17 10.13 15.39 11.92 17.85 10.95 16.24
+2:4 wanda 14.78 22.84 11.74 18.24 15.80 23.83 12.06 18.73
+[…]
+[…]
+[…]
+[…]
+Table 14: Wikitext2 and C4 perplexity (↓) for Qwen2.5 models, with a context length of 2048.
 OPT-1.3B OPT-2.7B OPT-6.7B OPT-13B
@@
 2:4 Wanda 28.27 28.54 21.17 22.84 15.90 18.99 15.55 16.18
-W4S20% 14.49 16.60 12.03 14.54 10.21 12.71 9.93 12.16
-W4S30% 16.06 18.44 13.23 15.95 10.94 13.64 10.37 12.85
-W4S40% 18.82 21.54 15.39 18.25 12.15 15.12 11.29 13.97
-W4S50% GQSA 21.32 24.90 17.52 20.81 13.44 16.94 12.16 15.57
+[…]
+[…]
+[…]
+[…]
 Table 15: Wikitext2 and C4 perplexity (↓) for OPT models, with a context length of 2048.
-Appendix H GQSA inference latency and memory consumption
+Appendix M GQSA inference latency and memory consumption
 Due to space constraints, detailed inference latency and model memory consumption are provided in Appendix Table 16. Overall, GQSA demonstrates exceptional performance across various settings.